# Well-posedness of a generalized Stokes operator on domains with cylindrical ends via layer-potentials

### 11 May 2026 v1

@@ Abstract.
on a domain with straight cylindrical ends $\Omega$ using the method of layer potentials on a larger manifold with straight cylindrical ends $M$, $\Omega\subset M$. The operator $\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}$ recovers the classical Stokes operator $\boldsymbol{\Xi}_{0,0}$ when the potentials $V$ and $V_{0}$ vanish. Under suitable positivity assumptions at infinity on $V$ and $V_{0}$, we prove that $\boldsymbol{\Xi}$ is Fredholm on $L^{2}(M;TM\oplus\mathbb{C})$. This allows us then to define the single- and double-layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$. Under further positivity assumptions at infinity, we prove that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are also Fredholm. Under slightly stronger assumptions on $V$ and $V_{0}$ (including positivity everywhere), we prove the invertibility of the operators $\boldsymbol{\Xi}$, $\boldsymbol{S}$, and $\frac{1}{2}+\boldsymbol{K}$. As it is well-known, the invertibility of these operators leads to well-posedness results for the associated (linear) Stokes boundary value problem with Dirichlet boundary conditions on domains with straight cylindrical ends. The proofs of these results required us to develop many other related tools. In particular, we develop an “algebra tool kit” to deal with limit and jump relations of layer potential operators, in general, and on manifolds with straight cylindrical ends, in particular. We do that first on $\mathbb{R}^{n}$, then on (possibly) non-compact manifolds, and then we deal with the specific case of manifolds with straight cylindrical ends. We use these results to study the limit and jump relations of the layer potential operators associated to our generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ on manifolds with cylindrical ends. We also develop Green formulas and energy estimates for our generalized Stokes operator $\boldsymbol{\Xi}$ on manifolds with straight cylindrical ends, which requires, in particular, a careful geometric study of the related differential operators, such as the deformation operator $\operatorname{Def}\,$. We also use suitable classes of pseudodifferential operators on manifolds with straight cylindrical ends that were studied in some previous papers of ours (including “The Stokes operator on manifolds with cylindrical ends,” J. Diff. Equations, 2024). For completeness, we include a review of the needed results on these pseudodifferential operators. As an application of all these results, we prove the well-posedness result for the Dirichlet problem for the generalized Navier-Stokes system with small data on a domain with cylindrical ends. We expect our results to have applications to the study of Navier-Stokes equations on domains with conical points.
@@ 1. Introduction
We prove the well-posedness of the Dirichlet boundary value problem for a generalized Stokes system (1.4) on a domain with straight cylindrical ends $\Omega$ using the method of layer potentials. In the process, we obtain many other results of independent interest on the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (1.3), on related differential operators, and on their associated layer potentials. We carefully develop the needed concepts and methods and prove in detail our results. Before stating in more detail our main results, let us introduce the setting of our paper.
1.1. The setting of our well-posedness results
We will work on manifolds and domains with straight cylindrical ends (see the next figure, Figure 1, as well as Figure 2 at the beginning of Section 6).
@@ 1.1. The setting of our well-posedness results
More precisely, by a manifold with straight cylindrical ends $M$ we will mean a non-compact Riemannian manifold without boundary (i.e., boundaryless) isometric to one of the form
@@
  • $(M_{0},g_{0})$ is a compact Riemannian manifold,
@@
(see, for instance, [63, 67, 68]). There is no loss of generality to assume that $M$ is connected, so we will do that throughout the paper, in order to simplify our statements. (Also, we may assume $R_{M}=0$, which we will often do.)
Second, $\Omega$ will be a smooth domain with straight cylindrical ends in $M$, in the sense that
@@
The generalized Stokes operators $\boldsymbol{\Xi}_{V,V_{0}}$ are defined as follows (see Section 2 for more details, as well as for the concepts not recalled here). Let
@@
be the deformation operator, (see formula (2.1) and the discussion around it for more details), and $\boldsymbol{L}:=2\operatorname{Def}^{*}\,\operatorname{Def}\,$ be the “Deformation Laplacian.” Let $\operatorname{End}(E)$ be the set of linear maps $E\to E$ (the endomorphisms of $E$). For suitable “potentials” $V\in\mathcal{C}^{\infty}(M;\operatorname{End}(TM))$ and $V_{0}\in\mathcal{C}^{\infty}(M)$, we consider the operator
@@
Let
@@
so that $\boldsymbol{u}$ is the “vector part” of $U$. We then consider the Dirichlet boundary value problem
@@ 1.2. Main results
Our main results are for manifolds with straight cylindrical ends $M$ and for compatible open subsets $\Omega\subset M$ (with boundary $\Gamma$, as above) or for closed manifolds (i.e., compact, smooth, boundaryless manifolds). The case of closed manifolds was treated in great detail in [69], so we will largely ignore this case in this paper (but we will recall the results for the closed case in Subsection 12.2 following [69]). Let us thus assume in this introduction that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a manifold with straight cylindrical ends (see 1.1 or Definition 3.11); this will be the case for the most part of our paper. Our first main results are to prove the invertibility on $\Gamma$ of the single layer and double layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$, associated to suitable generalizations $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ of the Stokes operator $\boldsymbol{\Xi}_{0,0}$, which will be defined shortly, where $V$ and $V_{0}$ are suitable non-negative functions, see Theorems 10.4 and 10.8. (We usually denote the identity operator $\mathbb{I}$ with 1, we thus write $\frac{1}{2}+\boldsymbol{K}$ instead of $\frac{1}{2}\mathbb{I}+\boldsymbol{K}$.) We also obtain, for suitable potentials $V$ and $V_{0}$ satisfying slightly weaker assumptions, that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are Fredholm.
@@
In order to define our layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$, we first prove that $\boldsymbol{\Xi}$ is Fredholm under suitable positivity assumptions at infinity for $V$ and $V_{0}$. We prove that it ($\boldsymbol{\Xi}$) is even invertible if the positivity conditions hold everywhere (Theorem 8.4). We use these results to define the single and double layer potential operators $\mathcal{S}_{\rm{ST}}$, $\boldsymbol{S}$, $\mathcal{D}_{\rm{ST}}$, and $\boldsymbol{K}$ associated to the Dirichlet boundary value problem for $\boldsymbol{\Xi}$. We prove that they satisfy the usual mapping and “jump” properties and that they have the usual principal symbols (Theorem 8.16; but in order to obtain this result, we develop an extensive theory: energy and representation formulas, geometric representations and formulas for our operators, a “took-kit” for proving jump relations, and many other results of independent interest). Then, using the jump relations and principal symbol properties of the layer potential operators, we prove that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are invertible. As is well known, this result implies the solvability of the Dirichlet problem (1.4), Theorem 1.1 next, which is the main result of this paper. We write $v\succ 0$ for a non-negative function that does not vanish identically on any connected component of its domain (see Definition 2.13). The space $\mathcal{C}^{\infty}_{\operatorname{inv}}(M)$ consists of functions (or sections of a vector bundle in case of $\mathcal{C}^{\infty}_{\operatorname{inv}}(M;E)$) that are translation invariant in a neighborhood of invinity (Definition 3.14)
@@
    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, and
@@
    if $\Omega_{0}$ is a connected component of $M^{\prime}$ contained in $\Omega^{\prime}$, then $\color[rgb]{1,0,1}\widehat{V}\not\equiv 0$ on $\Omega_{0}$.
@@
The proof of this theorem is provided in Subsection 11.1. At this time, a proof in general of this theorem using energy methods (i.e., the Riesz representation formula) seems out of reach. We also obtain applications of this well-posedness results to the Dirichlet-to-Neumann operator and to the solvability of the stationary Navier-Stokes system on domains with straight cylindrical ends of small dimension.
We have outlined here the most important results. However, in this paper, we obtain many other results of independent interest that lead to the main results. These other results are discussed in much more detail in the part on the contents of the paper, which is next.
@@ 1.3. Contents of the paper
Our paper consists of 12.1 sections (including an introduction), each divided into subsections. The assumptions of this paper (especially on the manifold $M$ on which we work) change from section to section. These assumptions are always stated at the beginning of that section and are reminded regularly, especially at the beginning of the subsections and in the statement of the most important results. (Note, however, that, beginning with Section 5, we work almost exclusively on manifolds with straight cylindrical ends.)
@@
The twelfth (and last) section is devoted to applications and extensions. As an application, we obtain an existence and uniqueness result for the Dirichlet problem for the generalized Navier-Stokes system (12.4) in a smooth domain $\Omega$ with straight cylindrical ends assuming $k\in\mathbb{Z}_{+}$, $\operatorname{dim}\,M=n<2(k+3)$, that the given data is in $H^{k}(\Omega;TM)\oplus H^{k+3/2}(\partial\Omega;TM)$ and is sufficiently small in a sense that is described in detail in Theorem 12.3. The proof is based on the well-posedness of the Dirichlet problem for the non-homogeneous generalized Stokes system, Theorem 11.4, and a fixed point theorem. In the last subsection of this section (and of this paper), we state the well-posedness of the Dirichlet problem for the generalized Stokes system on a closed manifold [69]. We also explain what modifications need to be brought to the statements of our results for manifolds with straight cylindrical ends so that they are valid also in the case of closed manifolds. These resuls were proved in [69] (for closed manifolds), so in this paper we content ourselves with reminding a few of these results.
@@ 1.4. Overview of the existing literature
The layer potentials method is a classical method for solving elliptic boundary value problems [32, 35, 37, 39, 45, 54, 55, 59, 63, 76, 87, 96, 117, 134]. It plays a crucial role in the analysis of elliptic boundary value problems in various function spaces, their main advantage consisting in the fact that they provide explicit formulas for the the solutions and that they provide a convenient way of extending the results to $L^{p}$-type Sobolev spaces. There are also results that can be proved at this time only using layer potential methods. An example is provided by our well-posedness theorem, Theorem 1.1. More on the history of the method of layer potentials can be found, for example, in [2, 43, 45, 54, 59, 76, 87, 115, 135].
@@
Let us now review some of the more specialized results in the subject, most of them from research papers devoted to the analysis of the Stokes and Navier-Stokes equations on domains, in Euclidean spaces, or on Riemannian manifolds. Fabes, Kenig and Verchota [43] initiated the study of boundary value problems for the Stokes system in Lipschitz domains of $\mathbb{R}^{n}$ with data in $L^{2}$-based Sobolev spaces and obtained various mapping properties for the corresponding layer potentials. They also obtained well-posedness results for the Dirichlet and Neumann problems in $L^{p}$-based spaces, with $p$ in a neighborhood of $2$, by combining Rellich type formulas with layer potential approaches. Mitrea and Wright [100] used layer potential methods and extended the results in [43] to a variety of $L^{p}$, Sobolev, Bessel potential, and Besov spaces. They also obtained well-posedness results for the main boundary value problems for the constant coefficient Stokes system in arbitrary Lipschitz domains in Euclidean spaces, together with the determination of the optimal range for $p$. The authors of [61] obtained mapping properties for the constant-coefficient Stokes and Brinkman layer potential operators in standard and weighted Sobolev spaces on $\mathbb{R}^{3}$. Shen [127] used a layer potential approach and obtained resolvent estimates in $L^{p}$ spaces for the Stokes operator in Lipschitz domains in the Euclidean spaces. Layer potential theoretical methods have been also combined with fixed point theorems in the analysis of boundary value problems for (linear and nonlinear) elliptic systems with nonlinear boundary conditions. For instance, the authors in [62] used a layer potential approach and the Leray-Schauder fixed point theorem to show the existence result in $L^{p}$, Sobolev, and Besov spaces for the constant-coefficient Stokes and Brinkman systems with nonlinear Neumann-transmission conditions. Girault and Sequeira [46] used a variational approach to prove the well-posedness of the exterior Dirichlet problem for the Stokes system on exterior Lipschitz domains in $\mathbb{R}^{n}$, $n=2,3$. Galdi [45, Chapters 5, 11, 12] used variational approaches in the analysis of steady Stokes and Navier-Stokes flows in exterior domains. Medková [90] used an integral equation method to analyze the Dirichlet problem for the Brinkman and Darcy-Forchheimer-Brinkman systems in Sobolev spaces on unbounded domains with compact Lipschitz boundary in $\mathbb{R}^{n}$. Exterior Dirichlet and transmission problems for anisotropic Stokes and Navier-Stokes systems with $L_{\infty}$ variable coefficients have been studied in [65] and [63, Chapters 5-7] using a variational approach and in [92]. Other boundary problems for the Stokes system in Sobolev spaces on Lipschitz domains via layer potential theoretical methods have also been studied in [88, 89, 91, 115, 118, 136].
@@
An important paper is Costabel’s contribution [35] that establishes many very useful properties of layer potentials using elementary methods. Chandler-Wilde et al. [31, 32] studied boundary integral equations on locally-dilation-invariant Lipschitz domains and fractal screens. Buffa, Hiptmair, Petersdorff, and Schwab [28] studied Maxwell transmission problems in Lipschitz domains by using boundary element methods. Labarca-Figueroa and Hiptmair [75] used coupled boundary and volume integral equations in the study of electromagnetic scattering. Singularly perturbed boundary value problems based on a functional analytic approach have been studied by M. Dalla Riva, M. Lanza de Cristoforis, and P. Musolino [37]. This approach, which also uses a layer potential analysis, is very useful in the study of various linear nonlinear elliptic problems. A nonvariational form of the acoustic single layer potential has been obtained by Lanza de Cristoforis [78]. A nonvariational form of the Neumann problem for Hölder continuous harmonic functions, based also on a layer potential analysis, has been obtained in [77]. Ando, Kang, Miyanishi, and Putinar have obtained Carleman factorizations of layer potentials [15]. The recent book [96] of D. Mitrea, I. Mitrea and M. Mitrea is a rigorous interplay between Harmonic Analysis, Geometric Measure Theory, Function Space Theory, and Partial Differential Equations, and provide a powerful tool in the analysis of boundary problems for complex coefficient elliptic systems in various geometric settings, including the class of Lipschitz domains. The theory of Fredholm and layer potential operators in Euclidean spaces and in Riemannian manifolds plays an important role in their book [96] (and also in our book [63]).

### 28 May 2026 v2

@@ Abstract.
on a domain with straight cylindrical ends $\Omega$ using the method of layer potentials on a larger manifold with straight cylindrical ends $M$, $\Omega\subset M$. The operator $\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}$ recovers the classical Stokes operator $\boldsymbol{\Xi}_{0,0}$ when the potentials $V$ and $V_{0}$ vanish. Under suitable positivity assumptions at infinity on $V$ and $V_{0}$, we prove that $\boldsymbol{\Xi}$ is Fredholm on $L^{2}(M;TM\oplus\mathbb{C})$. This allows us then to define the single- and double-layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$. Under further positivity assumptions at infinity, we prove that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are also Fredholm. Under slightly stronger assumptions on $V$ and $V_{0}$ (including non-negativity everywhere, i.e., $V,V_{0}\geq 0$), we prove the invertibility of the operators $\boldsymbol{\Xi}$, $\boldsymbol{S}$, and $\frac{1}{2}+\boldsymbol{K}$. As it is well-known, the invertibility of these operators leads to well-posedness results for the associated (linear) Stokes boundary value problem with Dirichlet boundary conditions on domains with straight cylindrical ends. The proofs of these results required us to develop many other related tools. In particular, we develop an “algebra tool kit” to deal with limit and jump relations of layer potential operators, in general, and on manifolds with straight cylindrical ends, in particular. We do that first on $\mathbb{R}^{n}$, then on (possibly) non-compact manifolds, and then we deal with the specific case of manifolds with straight cylindrical ends. We use these results to study the limit and jump relations of the layer potential operators associated to our generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ on manifolds with cylindrical ends. We also develop Green formulas and energy estimates for our generalized Stokes operator $\boldsymbol{\Xi}$ on manifolds with straight cylindrical ends, which requires, in particular, a careful geometric study of the related differential operators, such as the deformation operator $\operatorname{Def}\,$. We also use suitable classes of pseudodifferential operators on manifolds with straight cylindrical ends that were studied in some previous papers of ours (including “The Stokes operator on manifolds with cylindrical ends,” J. Diff. Equations, 2024). For completeness, we include a review of the needed results on these pseudodifferential operators. As an application of all these results, we prove the well-posedness result for the Dirichlet problem for the generalized Navier-Stokes system with small data on a domain with cylindrical ends. We expect our results to have applications to the study of Navier-Stokes equations on domains with conical points.
@@ 1. Introduction
We prove the well-posedness of the Dirichlet boundary value problem for a generalized Stokes system
[…]
for a domain with straight cylindrical ends $\Omega$ using the method of layer potentials. (Here $\nabla$ is the gradient, $\boldsymbol{L}$ is the deformation Laplacian, and $V$ and $V_{0}$ are potentials, see the next subsection. See also Equation (1.5) for a reformulation of this system using the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (1.4).) In the process, we obtain many other results of independent interest on the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$, on related differential operators, and on their associated layer potentials. We carefully develop the needed concepts and methods and prove in detail our results.
Before stating our main results, let us first introduce the setting of our paper.
1.1. The setting of our well-posedness results: Domains and manifolds with straight cylindrical ends
We will work on manifolds $M$ and domains $\Omega$ with straight cylindrical ends (see the next figure, Figure 1, as well as Figure 2 at the beginning of Section 6 for pictorial description of our setting).
@@ 1.1. The setting of our well-posedness results: Domains and manifolds with straight cylindrical ends
1.1.1. Manifolds with straight cylindrical ends
Intuitively, a manifold with straight cylindrical ends is one that, outside a bounded set, it is a (half-)infinite cylinder $M^{\prime}\times(-\infty,R]$ with a product metric, where $M^{\prime}$ is a closed manifold (i.e. a smooth, compact manifold without boundary).
More precisely, by a manifold with straight cylindrical ends $M$ we will mean a non-compact Riemannian manifold without boundary (i.e., “boundaryless”) isometric to one of the form
@@ 1.1.1. Manifolds with straight cylindrical ends
  • $(M_{0},g_{0})$ is a compact Riemannian manifold with boundary $\partial M_{0}$,
@@
(where we followed the conventions in [83, 87, 88]; but see also [23, 27, 29, 30, 54, 56, 94], where manifolds with cylindrical ends arise in the setting of “wave-guides”). There is no loss of generality to assume that $M$ is connected, so we will do that throughout the paper, in order to simplify our statements. (Also, we may assume $R_{M}=0$, which we will often do.)
1.1.2. Domains with straight cylindrical ends
In fact, we are more interested in the smooth domain with straight cylindrical ends $\Omega$, on which our boundary value problem (1.5) is formulated. Often $\Omega$ will be contained naturally in a manifold with straight cylindrical ends $M$ in a compatible way, in the sense that
@@ 1.1.2. Domains with straight cylindrical ends
1.1.3. The Dirichlet boundary value problem for the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$
Our main interest is, in fact, in the Dirichlet boundary value problem for a generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$, this problem is formulated below, see Equation (1.5). The generalized Stokes operators $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ are defined as follows. Let
@@
be the deformation operator (see formula (2.1) for the definition of $\operatorname{Def}\,$ and, in general, see Section 2 for more details). Let then $\boldsymbol{L}:=2\operatorname{Def}^{*}\,\operatorname{Def}\,$ be the “Deformation Laplacian.” Let $\operatorname{End}(E)$ be the set of linear maps $E\to E$ (the endomorphisms of $E$). For suitable “potentials” $V\in\mathcal{C}^{\infty}(M;\operatorname{End}(TM))$ and $V_{0}\in\mathcal{C}^{\infty}(M)$ that are translation invariant in a neighborhood of infinity (Definition 3.14), we consider the operator
@@
This operator extends to Sobolev spaces in distribution sense. Let
@@
so that $\boldsymbol{u}$ is the “vector part” of $U$. We then consider the Dirichlet boundary value problem, Equation (1.1), which, in this notation, becomes
@@ 1.2. Main results
Our main results are for manifolds with straight cylindrical ends $M$ and for compatible open subsets $\Omega\subset M$ with boundary $\Gamma:=\partial\Omega$, as above. Our main results also hold for closed manifolds (i.e., compact, smooth, boundaryless manifolds), see [89] (some of these results are recalled in Subsection 12.2). Let us thus assume in this introduction that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a manifold with straight cylindrical ends (see 1.1 or Definition 3.11); this will be the case for the most part of our paper.
Our first main results are to prove the invertibility on $\Gamma$ of the single layer and double layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$, associated to suitable generalizations $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ of the Stokes operator $\boldsymbol{\Xi}_{0,0}$, which will be defined shortly, where $V$ and $V_{0}$ are suitable non-negative potentials (as in 1.1), see Theorems 10.4 and 10.8. (We usually denote the identity operator $\mathbb{I}$ with 1, we thus write $\frac{1}{2}+\boldsymbol{K}$ instead of $\frac{1}{2}\mathbb{I}+\boldsymbol{K}$.) We also obtain, for suitable potentials $V$ and $V_{0}$ satisfying slightly weaker assumptions, that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are Fredholm.
@@
In order to define our layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$, we first prove that $\boldsymbol{\Xi}$ is Fredholm under suitable positivity assumptions at infinity for $V$ and $V_{0}$. We prove that it ($\boldsymbol{\Xi}$) is even invertible if the positivity conditions hold everywhere (Theorem 8.4). We use these results to define the single and double layer potential operators $\mathcal{S}_{\rm{ST}}$, $\boldsymbol{S}$, $\mathcal{D}_{\rm{ST}}$, and $\boldsymbol{K}$ associated to the Dirichlet boundary value problem for $\boldsymbol{\Xi}$. We prove that they satisfy the usual mapping and “jump” properties and that they have the usual principal symbols (Theorem 8.16; but in order to obtain this result, we develop an extensive theory: energy and representation formulas, geometric representations and formulas for our operators, the definition and suitable classes of integral operators, a “took-kit” for proving jump relations, and many other results of independent interest). Then, using the jump relations and principal symbol properties of the layer potential operators, we prove that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are invertible. As is well known, this result implies the solvability of the Dirichlet problem (1.5) (see Equation (1.5) for a more explicit form of this equation), Theorem 1.1 next, which is the main result of this paper. We write $v\succ 0$ for a non-negative function that does not vanish identically on any connected component of its domain (see Definition 2.13). The space $\mathcal{C}^{\infty}_{\operatorname{inv}}(M)$ consists of functions (or sections of a vector bundle in case of $\mathcal{C}^{\infty}_{\operatorname{inv}}(M;E)$) that are translation invariant in a neighborhood of infinity (Definition 3.14).
@@
    $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, and
@@
    if $\Omega_{0}$ is a connected component of $M^{\prime}$ contained in $\Omega^{\prime}$, then $\widehat{V}\not\equiv 0$ on $\Omega_{0}$.
@@
The proof of this theorem is provided in Subsection 11.1. At this time, a proof in general of this theorem using energy (or variational) methods (i.e., the Riesz representation formula) seems out of reach, in the sense that it does not follow from the Riesz representation formula using standard methods. We also obtain applications of this well-posedness results to the Dirichlet-to-Neumann operator and to the solvability of the stationary Navier-Stokes system on domains with straight cylindrical ends of small dimension.
We have outlined here the most important results. However, in this paper, we obtain many other results of independent interest that lead to the main results (they were mentioned briefly above). We also include a lot of background material, which is, hence not entirely new. However, our presentation is original and it gathers many results that are difficult to find in earlier papers. Moreover, most of our background material is presented in great detail, so that to make our main result more accessible (a notable exception is, though, the background material on pseudodifferential operators, for which many other well-developed works exist). These other results are discussed in much more detail in the part on the contents of the paper, which is next.
@@ 1.3. Contents of the paper
Our paper consists of 12 sections (including an introduction), each divided into subsections. The assumptions of this paper (especially on the manifold $M$ on which we work) change from section to section. These assumptions are always stated at the beginning of that section and are reminded regularly, especially at the beginning of the subsections and in the statement of the most important results. (Note, however, that, beginning with Section 5, we work almost exclusively on manifolds with straight cylindrical ends.) The first six sections are devoted mostly to background material, but the presentation is original (and, except for the part on pseudodifferential operators, quite detailed, including details and even results that are hard to find in the literature).
@@
The twelfth (and last) section is devoted to applications and extensions. As an application, we obtain an existence and uniqueness result for the Dirichlet problem for the generalized Navier-Stokes system (12.4) in a smooth domain $\Omega$ with straight cylindrical ends assuming $k\in\mathbb{Z}_{+}$, $\operatorname{dim}\,M=n<2(k+3)$, that the given data is in $H^{k}(\Omega;TM)\oplus H^{k+3/2}(\partial\Omega;TM)$ and is sufficiently small in a sense that is described in detail in Theorem 12.3. The proof is based on the well-posedness of the Dirichlet problem for the non-homogeneous generalized Stokes system, Theorem 11.4, and a fixed point theorem. In the last subsection of this section (and of this paper), we state the well-posedness of the Dirichlet problem for the generalized Stokes system on a closed manifold [89]. We also explain what modifications need to be brought to the statements of our results for manifolds with straight cylindrical ends so that they are valid also in the case of closed manifolds. These results were proved in [89] (for closed manifolds), so in this paper we content ourselves with reminding a few of these results.
@@ 1.4. Overview of the existing literature
The layer potentials method is a classical method for solving elliptic boundary value problems [39, 42, 47, 49, 59, 71, 72, 73, 79, 83, 96, 108, 118, 140, 158]. It plays a crucial role in the analysis of elliptic boundary value problems in various function spaces, their main advantage consisting in the fact that they provide explicit formulas for the the solutions and that they provide a convenient way of extending the results to $L^{p}$-type Sobolev spaces. There are also results that can be proved at this time only using layer potential methods. An example is provided by our well-posedness theorem, Theorem 1.1. More on the history of the method of layer potentials can be found, for example, in [2, 57, 59, 71, 74, 79, 96, 108, 138, 159]. See also [51, 60, 75].
@@
Let us now review some of the more specialized results in the subject, most of them from research papers devoted to the analysis of the Stokes and Navier-Stokes equations on domains, in Euclidean spaces, or on Riemannian manifolds. Fabes, Kenig and Verchota [57] initiated the study of boundary value problems for the Stokes system in Lipschitz domains of $\mathbb{R}^{n}$ with data in $L^{2}$-based Sobolev spaces and obtained various mapping properties for the corresponding layer potentials. They also obtained well-posedness results for the Dirichlet and Neumann problems in $L^{p}$-based spaces, with $p$ in a neighborhood of $2$, by combining Rellich type formulas with layer potential approaches. Mitrea and Wright [122] used layer potential methods and extended the results in [57] to a variety of $L^{p}$, Sobolev, Bessel potential, and Besov spaces. They also obtained well-posedness results for the main boundary value problems for the constant coefficient Stokes system in arbitrary Lipschitz domains in Euclidean spaces, together with the determination of the optimal range for $p$. The authors of [81] obtained mapping properties for the constant-coefficient Stokes and Brinkman layer potential operators in standard and weighted Sobolev spaces on $\mathbb{R}^{3}$. Shen [150] used a layer potential approach and obtained resolvent estimates in $L^{p}$ spaces for the Stokes operator in Lipschitz domains in the Euclidean spaces. Layer potential theoretical methods have been also combined with fixed point theorems in the analysis of boundary value problems for (linear and nonlinear) elliptic systems with nonlinear boundary conditions. For instance, the authors in [82] used a layer potential approach and the Leray-Schauder fixed point theorem to show the existence result in $L^{p}$, Sobolev, and Besov spaces for the constant-coefficient Stokes and Brinkman systems with nonlinear Neumann-transmission conditions. Girault and Sequeira [61] used a variational approach to prove the well-posedness of the exterior Dirichlet problem for the Stokes system on exterior Lipschitz domains in $\mathbb{R}^{n}$, $n=2,3$. Galdi [59, Chapters 5, 11, 12] used variational approaches in the analysis of steady Stokes and Navier-Stokes flows in exterior domains. Medková [111] used an integral equation method to analyze the Dirichlet problem for the Brinkman and Darcy-Forchheimer-Brinkman systems in Sobolev spaces on unbounded domains with compact Lipschitz boundary in $\mathbb{R}^{n}$. Exterior Dirichlet and transmission problems for anisotropic Stokes and Navier-Stokes systems with $L_{\infty}$ variable coefficients have been studied in [85] and [83, Chapters 5-7] using a variational approach and in [113]. Other boundary problems for the Stokes system in Sobolev spaces on Lipschitz domains via layer potential theoretical methods have also been studied in [109, 110, 112, 138, 141, 160].
@@
Chandler-Wilde et al. [38, 39] studied boundary integral equations on locally-dilation-invariant Lipschitz domains and fractal screens. Cakoni, Hsio, and Wendland have studied a mixed boundary value problem for the biharmonic equation using boundary integral equations. Haller-Dintelmann, R. and Jonsson, A. and Knees, D. and Rehberg have studied the regularity of a mixed, second order boundary value problem in divergence form [67]. Singularly perturbed boundary value problems based on a functional analytic approach have been studied by M. Dalla Riva, M. Lanza de Cristoforis, and P. Musolino [47]. This approach, which also uses a layer potential analysis, is very useful in the study of various linear nonlinear elliptic problems. A nonvariational form of the acoustic single layer potential has been obtained by Lanza de Cristoforis [98]. A nonvariational form of the Neumann problem for Hölder continuous harmonic functions, based also on a layer potential analysis, has been obtained in [97]. Ando, Kang, Miyanishi, and Putinar have obtained Carleman factorizations of layer potentials [18]. The recent book [118] of D. Mitrea, I. Mitrea and M. Mitrea is a rigorous interplay between Harmonic Analysis, Geometric Measure Theory, Function Space Theory, and Partial Differential Equations, and provide a powerful tool in the analysis of boundary problems for complex coefficient elliptic systems in various geometric settings, including the class of Lipschitz domains. The theory of Fredholm and layer potential operators in Euclidean spaces and in Riemannian manifolds plays an important role in their book [118] (and also in our book [83]). See also [34, 35, 43].
Layer potentials play an important role in applications, for instance in the study of numerical methods. In this direction, in an important paper, Costabel [42] has establishes many useful properties of layer potentials using elementary methods. Buffa, Hiptmair, Petersdorff, and Schwab [33] studied numerical methods for Maxwell’s transmission problems in Lipschitz domains by using boundary element methods. Labarca-Figueroa and Hiptmair [95] used coupled boundary and volume integral equations in the study of electromagnetic scattering. Steinbach and Wendland [152] have obtained resulsts on a Neumann series method for solving boundary integral equations. Costabel, Stephan, and Wendland have studied boundary integral equations of the first kind for the bi-Laplacian in a polygonal plane domain [45]. See also [19, 44, 53, 123, 124, 127].